# EvoRefuse: Evolutionary Prompt Optimization for Evaluation and Mitigation of LLM Over-Refusal to Pseudo-Malicious Instructions

### v2

@@ Abstract
Large language models (LLMs) frequently refuse to respond to pseudo-malicious instructions: semantically harmless input queries triggering unnecessary LLM refusals due to conservative safety alignment, significantly impairing user experience. Collecting such instructions is crucial for evaluating and mitigating over-refusals, but existing instruction curation methods, like manual creation or instruction rewriting, either lack scalability or fail to produce sufficiently diverse and effective refusal-inducing prompts. To address these limitations, we introduce EvoRefuse, a prompt optimization approach that generates diverse pseudo-malicious instructions consistently eliciting confident refusals across LLMs. EvoRefuse employs an evolutionary algorithm exploring the instruction space in more diverse directions than existing methods via mutation strategies and recombination, and iteratively evolves seed instructions to maximize evidence lower bound on LLM refusal probability. Using EvoRefuse, we create two novel datasets: EvoRefuse-Test, a benchmark of 582 pseudo-malicious instructions that outperforms the next-best benchmark with 140.41% higher average refusal triggering rate across 9 LLMs, 34.86% greater lexical diversity, and 40.03% improved LLM response confidence scores; and EvoRefuse-Align, which provides 3,000 pseudo-malicious instructions with responses for supervised and preference-based alignment training. LLaMA3.1-8B-Instruct supervisedly fine-tuned on EvoRefuse-Align achieves up to 14.31% fewer over-refusals than models trained on the second-best alignment dataset, without compromising safety. Our analysis with EvoRefuse-Test reveals models trigger over-refusals by overly focusing on sensitive keywords while ignoring broader context.
@@ 1 Introduction
Using EvoRefuse, we create two novel datasets: EvoRefuse-Test and EvoRefuse-Align. EvoRefuse-Test, a benchmark of 582 pseudo-malicious instructions for evaluating the LLM over-refusal problem, significantly outperforms the existing best alternatives by achieving 140.41% higher average refusal triggering rates across 9 LLMs, 34.86% greater lexical diversity, and 40.03% improved response confidence scores. EvoRefuse-Align offers 3,000 instances for safety alignment, comprising instruction-response pairs for Supervised Fine-Tuning (SFT) and preference pairs for Direct Preference Optimization (DPO) [9], with responses automatically generated by the LLM (i.e., Gpt-4o). Fine-tuning LLaMA3.1-8B-Instruct with EvoRefuse-Align reduces over-refusals by 14.31% using SFT and 40.04% using DPO, compared to next-best baseline datasets, while maintaining the original safety performance of LLaMA3.1-8B-Instruct.
@@
  • We construct two impactful datasets with EvoRefuse: EvoRefuse-Test, a benchmark achieving more challenging and robust LLM over-refusal evaluation (e.g., 140.41% higher refusal rate, 34.86% greater lexical diversity), and EvoRefuse-Align, enabling effective over-refusal mitigation (e.g., 14.31% fewer over-refusals) while preserving LLM safety.
@@ 3.2 Variational Approximation
The optimization problem then becomes finding $\boldsymbol{x}^{*}=\arg\max_{\boldsymbol{x}}\textbf{ELBO}(\boldsymbol{x})$. Our defined ELBO($\boldsymbol{x}$) in Eq. (4) specifically comprises the two expected terms reflecting response confidence and refuse probability. As in Appendix A.2, $\log p_{\boldsymbol{\theta}}(\boldsymbol{r}\mid\boldsymbol{x},\boldsymbol{s})$ increases monotonically with the ELBO($\boldsymbol{x}$).
@@ 3.3 Optimizing ELBO via EvoRefuse
[…]
@@
4:     Recombination: $S_{R}\leftarrow$ $N$ Safe results of $\mathcal{R}(x_{i},x_{j})$ with $x_{i},x_{j}\in X_{\text{top}}$
@@ 3.4 Pseudo-Malicious Instruction Dataset Construction
Using EvoRefuse, we constructed two novel datasets: EvoRefuse-Test and EvoRefuse-Align. For EvoRefuse-Test, designed to evaluate LLM over-refusals, we selected 800 diverse instructions from TRIDENT-Core [39], applied EvoRefuse for optimization, and after safety filtering with Gpt-4o, obtained 582 pseudo-malicious instructions that trigger high refusal rates across various LLMs. For EvoRefuse-Align, intended to mitigate over-refusals through alignment, we sampled 3,000 instructions from TRIDENT-Core and used Gpt-4o to generate paired helpful (chosen) and refusal (rejected) responses suitable for preference-based fine-tuning. Comprehensive implementation details are provided in Appendix A.6.
@@ 4 Experimental Setup
We used default decoding parameters for inference across all models. For alignment fine-tuning, we train LLaMA3.1-8B-Instruct for 3 epochs using LoRA [17] with a warmup ratio of 0.03 and learning rates of 2e-5 for SFT and 1e-5 for DPO. To ensure fair and consistent evaluation across models aligned with varying safety mechanisms, we standardized the safety criteria for all LLMs by prepending the same system prompt (see Appendix A.7) that instructs models to refuse unsafe instructions.
@@
In RQ1, we introduce two refusal metrics: Prefix Refusal Rate (PRR), based on matching predefined refusal prefixes [40], and Classifier Refusal Rate (CRR)1, computed using the classifier from fitness estimation. For Diversity, we use Mean segmental TTR (MSTTR), Hypergeometric distribution D (HDD)[41], and Measure of lexical textual diversity (MTLD)[42]; for Confidence, we use response Log-Probability (Log-Prob) and Long Text Perplexity (LongPPL) [43]; and for Safety, we annotate instructions manually into safe, debatable, and unsafe categories.
In RQ3, we evaluate the efficiency and stability of the optimization process using ELBO-based fitness scores and PRR. Full metric definitions are provided in the Appendix B.1.
@@ 5.1 EvoRefuse Enables a) Challenging and Robust Evaluation and b) Effective Mitigation
EvoRefuse-Test stands out as a more challenging benchmark for over-refusal evaluation, as evidenced by high refusal rates and high response confidence in Table 1. EvoRefuse-Test consistently induces the highest refusal rates across all evaluated LLMs. On average, it outperforms the second-best dataset SGTest by 140.41% across 9 models, with the most significant gain when testing LLaMA3.1-8B-Instruct (366.67%), likely because LLaMA3.1-8B-Instruct was the target LLM used in our pipeline. Importantly, EvoRefuse-Test also generalizes well to other LLMs, demonstrating that EvoRefuse discovers broadly effective over-refusal triggers rather than model-specific exploits. In contrast, datasets generated by baseline pipelines such as OR-Gen and PH-Gen yield significantly lower refusal rates, suggesting that evolutionary search more effectively explores instruction variants that reliably elicit refusals. Beyond refusal rates, EvoRefuse-Test also induces refusals with higher confidence. As shown in Table 2, it yields the highest average response log-probability and the lowest LongPPL among all benchmarks. Compared to the second-best dataset, this represents a 40.03% increase in log-probability and a 3.45% reduction in LongPPL.
@@
HITest 0.43 0.63 26.05 -77.91 1.61 0.92 0.08 0.00
OKTest 0.46 0.79 68.63 -86.06 1.12 0.91 0.06 0.03
OR-Bench 0.47 0.85 137.65 -93.45 1.26 0.88 0.10 0.02
OR-Gen 0.47 0.86 141.18 -99.12 1.18 0.92 0.07 0.01
PHTest 0.48 0.85 106.14 -94.60 1.16 0.89 0.07 0.04
PH-Gen 0.48 0.85 134.84 -103.08 1.15 0.89 0.09 0.02
SGTest 0.48 0.81 57.00 -83.67 1.28 0.93 0.05 0.02
XSTest 0.36 0.71 39.95 -72.62 1.34 0.97 0.03 0.00
EvoRefuse-Test 0.54 0.87 152.52 -43.55 1.12 0.94 0.06 0.00
Table 2: Evaluation of diversity, confidence, and safety on EvoRefuse-Test and baselines.
@@
+ Few Shots 0.98 0.94 0.96 0.95 0.57 0.55 0.14 0.13 0.14 0.14 0.73 0.75
+ DRO 0.96 0.94 0.95 0.93 0.58 0.55 0.10 0.10 0.15 0.17 0.72 0.72
@@
Table 3 compares strategies for mitigating over-refusals. Fine-tuning solely on TRIDENT-Core improves safety but significantly increases over-refusals. In contrast, both SFT and DPO using pseudo-malicious instructions reduce over-refusals while maintaining safety. Fine-tuning with EvoRefuse-Align achieves substantial improvement, reducing over-refusal rates by 14.31% compared to the best fine-tuning baseline and 14.76% compared to the best prompt-based method. Applying DPO with EvoRefuse-Align yields stronger results, reducing over-refusals by 40.04% with only a modest safety trade-off (3.87% decrease). These results demonstrate EvoRefuse-Align’s effectiveness in supporting alignment that balances safety preservation with over-refusal reduction.
@@ 6 Conclusion
In this work, we introduce EvoRefuse, an automated prompt optimization framework that generates diverse pseudo-malicious instructions that effectively induce mistaken refusal behavior in LLMs. By leveraging evolutionary search with a fitness objective derived from variational approximation, EvoRefuse produces two datasets: EvoRefuse-Test (582 examples), a robust and challenging refusal evaluation benchmark that elicits 140.41% higher average refusal rates across 9 LLMs, with 40.03% higher response confidence and 34.86% greater lexical diversity than the best existing benchmark; and EvoRefuse-Align (3,000 examples), an effective alignment dataset for refusal mitigation. Fine-tuning LLaMA3.1-8B-Instruct on EvoRefuse-Align reduces over-refusals by 14.31% under SFT and 40.04% under DPO while maintaining LLM safety. Analysis with EvoRefuse-Test reveals that LLMs trigger over-refusals by overly focusing on salient textual cues while ignoring broader linguistic context.

### v3

@@ Abstract
Large language models (LLMs) frequently refuse to respond to pseudo-malicious instructions: semantically harmless input queries triggering unnecessary LLM refusals due to conservative safety alignment, significantly impairing user experience. Collecting such instructions is crucial for evaluating and mitigating over-refusals, but existing instruction curation methods, like manual creation or instruction rewriting, either lack scalability or fail to produce sufficiently diverse and effective refusal-inducing prompts. To address these limitations, we introduce EvoRefuse, a prompt optimization approach that generates diverse pseudo-malicious instructions consistently eliciting confident refusals across LLMs. EvoRefuse employs an evolutionary algorithm exploring the instruction space in more diverse directions than existing methods via mutation strategies and recombination, and iteratively evolves seed instructions to maximize evidence lower bound on LLM refusal probability. Using EvoRefuse, we create two novel datasets: EvoRefuse-Test, a benchmark of 582 pseudo-malicious instructions that outperforms the next-best benchmark with 85.34% higher average refusal triggering rate across 9 LLMs without a safety-prior system prompt, 34.86% greater lexical diversity, and 40.03% improved LLM response confidence scores; and EvoRefuse-Align, which provides 3,000 pseudo-malicious instructions with responses for supervised and preference-based alignment training. With supervised fine-tuning on EvoRefuse-Align, LLaMA3.1-8B-Instruct achieves up to 29.85% fewer over-refusals than models trained on the second-best alignment dataset, without compromising safety. Our analysis with EvoRefuse-Test reveals models trigger over-refusals by overly focusing on sensitive keywords while ignoring broader context. Our code and datasets are available at https://github.com/FishT0ucher/EVOREFUSE.
**footnotetext: Corresponding authors.footnotetext: Senior author; led the research.
@@ 1 Introduction
Using EvoRefuse, we create two datasets: EvoRefuse-Test and EvoRefuse-Align. EvoRefuse-Test is a benchmark of 582 pseudo-malicious instructions for evaluating LLM over-refusal. It outperforms the strongest prior benchmark with 85.34% higher average refusal rates across 9 LLMs. When a safety-prior system prompt that instructs models to refuse unsafe instructions is enabled, this margin increases to 140.41%. EvoRefuse-Test also yields 34.86% greater lexical diversity and 40.03% higher response confidence. EvoRefuse-Align provides 3,000 instances for safety alignment, including instruction-response pairs for Supervised Fine-Tuning (SFT) and preference pairs for Direct Preference Optimization (DPO) [9], with responses generated by Gpt-4o. Fine-tuning LLaMA3.1-8B-Instruct on EvoRefuse-Align reduces over-refusals by 29.85% with SFT and 45.96% with DPO. When we enable the safety-prior system prompt that encourages refusals, the mitigation effect is smaller: over-refusals decrease by 14.31% (SFT) and 40.04% (DPO). In both settings, LLaMA3.1-8B-Instruct preserves its overall safety.
@@
  • We construct two impactful datasets with EvoRefuse: EvoRefuse-Test, a benchmark achieving more challenging and robust LLM over-refusal evaluation (e.g., 85.34% higher refusal rate, 34.86% greater lexical diversity), and EvoRefuse-Align, enabling effective over-refusal mitigation (e.g., 29.85% fewer over-refusals) while preserving LLM safety.
@@ 3.2 Variational Approximation
The optimization problem becomes finding $\boldsymbol{x}^{*}=\arg\max_{\boldsymbol{x}}\textbf{ELBO}(\boldsymbol{x})$. Our ELBO in Eq. (4) comprises two expected terms reflecting response confidence and refusal probability, and (under a fixed decoding scheme) treats the decoding entropy as an approximately constant offset. $\textbf{ELBO}(\boldsymbol{x})$ is a lower bound on $\log p_{\boldsymbol{\theta}}(\boldsymbol{r}\mid\boldsymbol{x},\boldsymbol{s})$. As clarified in Appendix A.2, increasing $\textbf{ELBO}(\boldsymbol{x})$ therefore improves this bound and typically correlates with larger $\log p_{\boldsymbol{\theta}}(\boldsymbol{r}\mid\boldsymbol{x},\boldsymbol{s})$, but it is not order-preserving in general. The true likelihood $\log p_{\boldsymbol{\theta}}(\boldsymbol{r}\mid\boldsymbol{x},\boldsymbol{s})$ may still fluctuate even when $\textbf{ELBO}(\boldsymbol{x})$ increases.
@@ 3.3 Optimizing ELBO via EvoRefuse
0: Seed instruction $x^{0}$, number of iterations II, number of recombinations NN, number of recombination candidates $L$, fitness evaluation function $\mathcal{F}(\cdot)$, collection of mutators $\mathcal{\textbf{M}}=$ $\mathcal{M_{\text{1}}}(\cdot),\mathcal{M_{\text{2}}}(\cdot),\dots$}, recombinator $\mathcal{R}(\cdot)$, safety classifier $\mathcal{J}(\cdot)$, cooling coefficient $\beta$, initial temperature $\tau_{0}$, final temperature $\tau_{f}$.
@@
4:  Recombination: $S_{R}\leftarrow$ NN Safe results of $\mathcal{R}(x_{i},x_{j})$ with $x_{i},x_{j}\in X_{\text{top}}$
@@
Under a fixed decoding scheme, the entropy term associated with the ELBO is treated as an approximately constant offset and is therefore omitted from $\mathcal{F}$. Thus, $\mathcal{F}(\boldsymbol{x})$ provides a Monte Carlo estimate of the practical surrogate in Eq. (4). Maximizing $\mathcal{F}(\boldsymbol{x})$ heuristically increases the practical ELBO and, by the ELBO identity, tends to improve the lower bound on $\log p_{\boldsymbol{\theta}}(\boldsymbol{r}\mid\boldsymbol{x},\boldsymbol{s})$, though it does not guarantee improvement of the true objective at every step.
@@ 3.4 Pseudo-Malicious Instruction Dataset Construction
Using EvoRefuse, we constructed two novel datasets: EvoRefuse-Test and EvoRefuse-Align. For EvoRefuse-Test, designed to evaluate LLM over-refusals, we selected 800 diverse instructions from TRIDENT-Core [39], applied EvoRefuse for optimization, and after safety filtering with Gpt-4o, obtained 582 pseudo-malicious instructions that trigger high refusal rates across various LLMs. For EvoRefuse-Align, intended to mitigate over-refusals through alignment, we sampled 3,000 instructions from TRIDENT-Core and used Gpt-4o to generate paired helpful (chosen) and refusal (rejected) responses suitable for preference-based fine-tuning. Comprehensive implementation details are provided in Appendix A.6, while the success rates of different mutation strategies and the results obtained with alternative LLMs as mutators are presented in the Appendix B.4 and B.6.
@@ 4 Experimental Setup
We keep other inference parameters at default values across models. For RQ1, we set the temperature to $0.01$ without a safety-prior system prompt and to $1.0$ with the safety-prior system prompt. For RQ2 and RQ3, we apply the safety-prior system prompt with temperature =1.0=1.0 to elicit more refusals, producing clearer signals for attribution analysis and more stable fitness/PRR trajectories during optimization. For alignment, we fine-tune LLaMA3.1-8B-Instruct for 5 epochs with LoRA [17] (warmup ratio 0.03; learning rate 2e-5 for SFT and 1e-5 for DPO). The system prompt is provided in Appendix B.5.
@@
In RQ1, we introduce two refusal metrics: Prefix Refusal Rate (PRR), based on matching predefined refusal prefixes [40], and Classifier Refusal Rate (CRR), which is computed using GPT as a classifier with the corresponding prompt provided in the appendix A.7. For Diversity, we use Mean segmental TTR (MSTTR), Hypergeometric distribution D (HDD)[41], and Measure of lexical textual diversity (MTLD)[42]; for Confidence, we use response Log-Probability (Log-Prob) and Long Text Perplexity (LongPPL) [43]; and for Safety, three NLP practitioners manually annotated the instructions into safe, debatable, and unsafe categories. In RQ2, we perform attribution analysis using gradient-based weights[44] and information flow[6] to visualize the contribution of different tokens to refusal prediction. In RQ3, we evaluate the efficiency and stability of the optimization process using ELBO-based fitness scores and PRR. Full metric definitions are provided in the Appendix B.1.
@@
Table 1: Evaluation refusal rates of LLMs on EvoRefuse-Test and baselines using PRR.
@@
HITest 0.08 0.12 0.04 0.00 0.00 0.04 0.08 0.04 0.20
OKTest 0.09 0.06 0.01 0.05 0.07 0.06 0.08 0.16 0.40
OR-Bench 0.14 0.15 0.05 0.04 0.07 0.09 0.27 0.06 0.18
OR-Gen 0.16 0.08 0.06 0.04 0.10 0.16 0.38 0.12 0.19
PHTest 0.10 0.19 0.08 0.09 0.03 0.10 0.12 0.09 0.31
PH-Gen 0.19 0.14 0.07 0.11 0.11 0.19 0.45 0.16 0.28
SGTest 0.18 0.14 0.14 0.00 0.05 0.09 0.12 0.14 0.32
XSTest 0.05 0.11 0.13 0.00 0.05 0.08 0.07 0.08 0.19
EvoRefuse-Test 0.24 0.26 0.65 0.12 0.25 0.27 0.38 0.24 0.74
@@ 5.1 EvoRefuse Enables a) Challenging and Robust Evaluation and b) Effective Mitigation
EvoRefuse-Test stands out as a more challenging benchmark for over-refusal evaluation, as evidenced by high refusal rates and high response confidence in Table 1 and Table 2. EvoRefuse-Test consistently induces the highest refusal rates across nearly all evaluated LLMs. On average, without a safety-prior system prompt, EvoRefuse-Test surpasses the strongest prior dataset PH-Gen by 85.34%. The largest gain is observed on LLaMA3.1-8B-Instruct (364.29%), likely because LLaMA3.1-8B-Instruct is the target LLM used in our pipeline. With a safety-prior system prompt enabled, EvoRefuse-Test further outperforms the next-best dataset SGTest by 140.41% across nine models, as the prompt amplifies refusal-prone behavior (Appendix B.5). Importantly, EvoRefuse-Test generalizes well beyond the target model, indicating that EvoRefuse discovers broadly effective over-refusal triggers rather than model-specific exploits. In contrast, datasets generated by baseline pipelines such as OR-Gen and PH-Gen yield significantly lower refusal rates, suggesting that evolutionary search more effectively explores instruction variants that reliably elicit refusals. Beyond refusal rates, EvoRefuse-Test also induces refusals with higher confidence. As shown in Table 2, it yields the highest average response log-probability and the lowest LongPPL among all benchmarks. Compared to the second-best dataset, this represents a 40.03% increase in log-probability and a 3.45% reduction in LongPPL.
Table 2: Evaluation of diversity, confidence, and safety on EvoRefuse-Test and baselines. $\pm$” shows the range across annotators.
@@
[…]
@@
LLaMA-3.1-Chat 0.94 0.95 0.94 0.95 0.53 0.60 0.11 0.10 0.14 0.15 0.65 0.66
+ Few Shots 0.97 0.97 0.99 0.99 0.53 0.56 0.12 0.12 0.21 0.22 0.48 0.49
+ DRO 1.00 1.00 0.98 0.99 0.64 0.63 0.14 0.15 0.14 0.14 0.56 0.53
+ TRIDENT-Core (SFT) 1.00 1.00 1.00 1.00 0.81 0.81 0.47 0.55 0.45 0.54 0.93 0.98
+ OR-Bench (SFT) 1.00 1.00 0.98 0.98 0.70 0.70 0.10 0.10 0.14 0.17 0.45 0.44
+ PHTest (SFT) 1.00 1.00 0.97 0.98 0.71 0.73 0.09 0.10 0.11 0.14 0.39 0.41
+ PromptAgent (SFT) 0.99 1.00 0.98 0.99 0.72 0.74 0.09 0.09 0.10 0.12 0.43 0.48
+ EvoRefuse-Align (SFT) 1.00 0.99 0.96 0.96 0.74 0.74 0.06 0.07 0.08 0.09 0.32 0.25
+ EvoRefuse-Align (DPO) 0.97 0.99 0.89 0.95 0.66 0.67 0.02 0.05 0.01 0.06 0.30 0.23
@@
Table 3 compares strategies for mitigating over-refusals. Fine-tuning solely on TRIDENT-Core improves safety but significantly increases over-refusals. In contrast, both SFT and DPO using pseudo-malicious instructions reduce over-refusals while maintaining safety. Fine-tuning with EvoRefuse-Align achieves substantial improvement, reducing over-refusal rates by 29.85% relative to the best fine-tuning baseline and by 46.95% relative to the best prompt-based method. Applying DPO with EvoRefuse-Align yields stronger results, reducing over-refusals by 45.96% with only a modest safety trade-off (4.82% decrease). With the safety-prior system prompt enabled, EvoRefuse-Align still reduces over-refusals by 14.31% (SFT) and 40.04% (DPO), while maintaining competitive overall safety (Appendix B.5). These results demonstrate EvoRefuse-Align’s effectiveness in supporting alignment that balances safety preservation with over-refusal reduction.
@@ 6 Conclusion
In this work, we introduce EvoRefuse, an automated prompt optimization framework that generates diverse pseudo-malicious instructions that effectively induce mistaken refusal behavior in LLMs. By leveraging evolutionary search with a fitness objective derived from variational approximation, EvoRefuse produces two datasets: EvoRefuse-Test (582 examples), a robust and challenging refusal evaluation benchmark that elicits 85.34% higher average refusal rates across 9 LLMs without a safety-prior system prompt, with 40.03% higher response confidence and 34.86% greater lexical diversity than the best existing benchmark; and EvoRefuse-Align (3,000 examples), an effective alignment dataset for refusal mitigation. Fine-tuning LLaMA3.1-8B-Instruct on EvoRefuse-Align reduces over-refusals by 29.85% under SFT and 45.96% under DPO while maintaining LLM safety. Analysis with EvoRefuse-Test reveals that LLMs trigger over-refusals by overly focusing on salient textual cues while ignoring broader linguistic context.
Author Contributions
Xiaorui Wu (Student Author) co-led conceptualization and methodology, led software development and experimental investigation, conducted formal analysis, wrote the original draft, and supported manuscript revision. Zhuang Li (Senior Author) co-led conceptualization and methodology, provided overall supervision and project direction, led project administration and manuscript revision, and contributed to validation and formal analysis. Fei Li contributed to methodology, supervision, and manuscript revision. Xiaofeng Mao contributed to methodology and manuscript revision. Chong Teng contributed to manuscript revision. Donghong Ji (Corresponding Author) contributed to project administration, resources, and manuscript revision. Xin Zhang (Corresponding Author) contributed to project administration, resources, and conceptualization. Yuxiang Peng and Li Zheng contributed to experimental investigation. All authors approved the final manuscript.
Acknowledgments
This work was supported by Ant Group and Wuhan University (the School of Cyber Science and Engineering) Joint Research Program on Large Language Model Safety Alignment (COOP246P255333), and the National Natural Science Foundation of China (No. 62176187).